# Autonomous Horizon-based Asteroid Navigation With Observability-constrained Maneuvers

### v1

@@ 6 Numerical Results
Following is a table that summarizes all the relevant penalty weights for both scenarios:
@@
Parameter Value
$\omega_{1}$ $1$
$k_{1}$ $1$
$\omega_{2}$ $1$
$k_{2}$ $1$
$\omega_{3.1}$ $10$
$\omega_{3.2}$ $100$
$k_{3}$ $1$
The penalty weight parameter subscript indicates the corresponding path constraint the parameter is relevant to according to the definitions in Section 4.1. All parameters have identical values except the cone weight, for which $\omega_{3.1}$ is used for the Asteroid Approach Targeting controller and $\omega_{3.2}$ is used for the Orbit Transfer controller.
@@ 6.1 Orbit Maintenance Scenario
To demonstrate the effectiveness of the observability-constrained controller, we develop a test scenario where a spacecraft starts with an arbitrary initial condition on a trajectory to flyby passing behind the dark side of Bennu. The target is a circular orbit with an inclination just above 30 degrees. This avoids the dark-side cone but provides a useful edge case to test where the spacecraft may often travel through the poor observability region in its controlled trajectory without an observability penalty. The initial conditions in cartesian terms are ${}^{H}\bm{r}=[1.0214,0,-2.0429]^{\top}$ km and ${}^{H}\bm{v}=[40.493,40.493,40.493]^{\top}$ mm/s. The gain found to provide the best performance in terms of stability and rate of convergence is as follows:
@@
$P_{0}=\mathrm{diag}(3.2761,3.2761,3.2761,8.544\times 10^{-17},8.544\times 10^{-17},8.544\times 10^{-17})$ (40)
To compare results, we test the Lyapunov controller without the observability improving penalty functions as shown in Fig. 9. The figure depicts the spacecraft’s true trajectory, OpNav measurement, EKF position estimate and a visualization of the keep-out cone. As can be seen in the results for the controller without an observability penalty, the spacecraft travels through the poor observability region near the asteroid’s dark side. This leads to a faulty measurement characterized by the spike in the OpNav measurement history, which causes the true and EKF-estimated trajectories to diverge.
We now compare this performance to that of the observability-constrained controller in Fig. 10. In the observability-constrained case, the controller successfully avoids the poor observability region, as can be seen by consistently matching the OpNav measurement and EKF prediction to the true trajectory at all times. An additional test was executed with a [2.5 1 1] ellipsoid using the rotation parameters of Bennu and an approximation from its known mean radius.
Camera snapshots of each case with the detected horizon points highlighted in red can be seen in Fig. 11. It displays the image history from the spacecraft camera perspective and allows us to make physical inferences about the controller’s performance. For the leftmost case without the observability-constrained control, the spacecraft does not avoid the keep-out cone which causes the lit limb to go out of view in the fourth and fifth images. Thus, the number of detected horizon points significantly decreases, causing the spacecraft to diverge. The limb comes back into view later, but is at a larger distance due to the divergence, making the measurement less effective. The middle case displays the same example but with the observability-constrained control, with which we see that the lit limb is kept within view throughout the trajectory. The rightmost case displays the view of the [2.5 1 1] ellipsoid using observability-constrained control to demonstrate the controller’s consistency in performance across varying shapes of celestial bodies. Although the size of the lit limb varies considerably due to the body’s rotation and viewing the ellipsoidal geometry from different angles, it maintains the lit limb within view by avoiding the poor observability zone and therefore produces more reliable measurements. Both the middle and rightmost cases avoid the divergence and disappearance of lit limb encountered in the leftmost case, supporting the benefit of the observability-constrained controller.
@@
Figure 9: True (blue), OpNav measurement (green), EKF-estimated (red) Trajectories and keep-out cone (shaded red) without Observability-constrained Penalty Lyapunov control around Bennu shown in 4 views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@
Figure 10: True (blue), OpNav measurement (green), EKF-estimated (red) Trajectories and keep-out cone (shaded red) with Observability-constrained Penalty Lyapunov control around Bennu shown in 4 views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@
Figure 11: Sample Image History with Detected Horizon Points from Fig. 10 for Case Without Observability-constrained control (Left), and With Observability-constrained control using Spherical Primary (Middle), and [2.5 1 1] Ellipsoid Primary (Right) Ordered Top-Bottom, then Left-Right.
@@ 6.1.1 Monte Carlo Analysis
To demonstrate the algorithm’s robustness, a Monte Carlo analysis is conducted to analyze its success rate statistically. A realistic initial standard deviation is defined for the position error across each direction ($x,y\text{ and }z$ in the Hill frame) of $\sigma=30$ meters, approximately 10% of Bennu’s radius. This is used to apply an initial error to the true initial state to generate the EKF’s initial known state. The Monte Carlo is run for 200 iterations. To start, we test the version without observability constraints to understand the severity of failures in this scenario. As per the results, the Lyapunov controller fails in almost every case without the observability constraint, as it passes through the poor observability region in every test. It succeeds in 8 out of 200 iterations for a success rate of 4%. The EKF and true dynamics diverge completely and lead to the spacecraft either crashing into the asteroid or traveling out of the system. Now, we compare this with the observability-constrained case. The observability-constrained Lyapunov controller shows better performance and success rate of the spacecraft reaching its desired orbit. It succeeds in 196 out of 200 iterations for a success rate of 98%. This is a major performance improvement and supports the utility of the observability-constrained controller. The ellipsoid case succeeds 161 out of 200 iterations for a success rate of 80.5%. The performance degrades with nonspherical objects but maintains an improvement over the case without an observability constraint. All Monte Carlo results are shown in Fig. 12.
@@
Figure 12: Monte Carlo Simulation of 200 True Trajectories using Basic Lyapunov controller (Left), Observability-constrained Lyapunov controller with Spherical (Middle) and [2.5 1 1] Ellipsoid Primary (Right). Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@ 6.2 Approach and Circularization Scenario
Beyond the orbit maintenance scenario, it is important to demonstrate the effectiveness of this controller in a greater variety of situations, including those starting further away from the asteroid. Therefore, we test this algorithm in a specific test case where we approach the asteroid from just within the effective range of horizon-based OpNav determined in Section 5.1. The controller aims to circularize around the asteroid and executes maneuvers accordigly. The initial conditions in cartesian terms are ${}^{H}\bm{r}=[0,0,-4.5964]^{\top}$ km and ${}^{H}\bm{v}=[-4.8927,0,2.4464]^{\top}$ mm/s. The same initial covariance is used as in Section 6.1. The gain found to return the best performance for the controller in this particular test case is:
@@
The penalty function parameters chosen are identical to those from Section 6.1, except the weight for the cone constraint is increased to $\omega_{3}$ = 100. This leads to better performance in avoiding the poor observability region at the larger scale of this scenario.
@@
We repeat a comparison between the regular (Fig. 13) and observability-constrained (Fig. 14) controllers. The figures depict the spacecraft’s true trajectory, OpNav measurement, EKF position estimate and a visualization of the keep-out cone. The result matches the outcome from Section 6.1 with the observability-constrained controller avoiding the keep-out cone and successfully circularizing around the asteroid. The regular controller’s true and EKF-estimated trajectories diverge since it enters the keep-out cone and produces faulty measurements. An additional test with the observability-constrained controller was executed with a [2.5 1 1] ellipsoid using the rotation parameters of Bennu and an approximation from its known mean radius.
@@
Camera snapshots of each case with the detected horizon points highlighted in red can be seen in Fig. 15. The observations are mostly similar to those from Section 6.1. The difference is that the lit limb is much smaller at the start of the simulation and grows as the spacecraft approaches the asteroid. The measurement and controller perform well under such variable conditions.
@@
The maximum fuel consumption to execute the control used during this trajectory was calculated to be approximately 0.1 kg over 5 days, which is a reasonable quantity within its fuel reserves. This calculation was done using the same method from Section 6.1.
@@
Figure 13: True (blue), OpNav measurement (green), EKF-estimated (red) Trajectories and keep-out cone (shaded red) without Observability-constrained Penalty Lyapunov control around Bennu shown in 4 views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@
Figure 14: True (blue), OpNav measurement (green), EKF-estimated (red) Trajectories and keep-out cone (shaded red) with Observability-constrained Penalty Lyapunov control around Bennu shown in 4 views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated.
@@
Figure 15: Sample Image History with Detected Horizon Points from Fig. 13 for Case Without Observability-constrained control (Left), and Fig. 14 for Cases With Observability-constrained control using Spherical Primary (Middle), and [2.5 1 1] Ellipsoid Primary (Right) Ordered Top-Bottom, then Left-Right.

### v2

@@ 5 Numerical Results
In the context of this paper, we define divergence as any case where the spacecraft violates the observability cone constraint or minimum and maximum distance safety constraints, which we consider as failure cases. We use this definition of divergence as the primary performance metric because the goal of our controller contribution is to ensure a spacecraft can safely and robustly attain a desired orbit around a primary small body without colliding into or leaving orbit around the small body, which it accomplishes and improves upon conventional control by adding observability constraints that ensure position measurements are persistently reliable along its controlled trajectory.
Divergence happens because of a combination of factors, starting with an extended period of erroneous measurements in a poor observability region due to violation of the cone constraint. These erroneous measurements cause the EKF’s state estimate to drift from the truth trajectory, which in turn leads to incorrect control input that exacerbates this deviation between truth and EKF. Thus, this leads to failure cases where the spacecraft violates the safety constraints, thereby colliding with the primary small body or exiting orbit around it. This is because the EKF estimate converges to an incorrect state knowledge matching the desired target orbit while the truth diverges with no corrective control, or erroneous control leads to the spacecraft persistently stuck in the poor observability region. Particularly in cases where the spacecraft leaves orbit around the small body, the OPNAV measurement can no longer compensate even after exiting the poor observability region, as its covariance expands at higher distances.
Table˜3 summarizes all the relevant penalty weights for both scenarios:
@@
Parameter Value Description
$\omega_{1}$ 11 Weight for min-radius constraint
$k_{1}$ 11 Sharpness for min-radius constraint
$\omega_{2}$ 11 Weight for max-radius constraint
$k_{2}$ 11 Sharpness for max-radius constraint
$\omega_{3.1}$ $10$ Weight for cone constraint (Controller 1)
$\omega_{3.2}$ $100$ Weight for cone constraint (Controller 2)
$k_{3}$ 11 Sharpness for cone constraint
The penalty weight parameter subscript indicates the corresponding path constraint the parameter is relevant to according to the definitions in Section˜3.1. All parameters have identical values except the cone weight, for which $\omega_{3.1}$ is used for the Small Body Approach Targeting controller and $\omega_{3.2}$ is used for the Orbit Transfer controller.
We reiterate that for all figures depicting distance norm error from either the OpNav measurement or EKF estimate in this paper, we use the square root of the maximum eigenvalue of the position covariance ($3\sqrt{\lambda_{max}(P_{\bm{r}})}$) to define the error bounds.
@@ 5.1 Orbit Maintenance Scenario
To demonstrate the effectiveness of the observability-constrained controller, we develop a test scenario where a spacecraft starts with an arbitrary initial condition on a trajectory to flyby passing behind the dark side of the small body. The target is a circular orbit with an inclination just above 30 degrees. This avoids the dark-side cone but provides a useful edge case to test where the spacecraft may often travel through the poor observability region in its controlled trajectory without an observability penalty. The initial conditions in Cartesian terms are $^{H}\bm{r}=[1.0214,0,-2.0429]^{\top}$ km and $^{H}\bm{v}=[40.493,40.493,40.493]^{\top}$ mm/s. Using the tuning guidelines outlined in appendix˜D, the gain found to provide the best performance in terms of stability and rate of convergence is as follows:
@@
To compare results, we test the Lyapunov controller without the observability improving penalty functions as shown in Fig.˜4. The figure depicts the spacecraft’s true trajectory, OpNav measurement, EKF position estimate and a visualization of the keep-out cone. As can be seen in the results for the controller without an observability penalty, the spacecraft travels through the poor observability region near the small body’s dark side. This leads to a faulty measurement, which causes the true and EKF-estimated trajectories to diverge.
We now compare this performance to that of the observability-constrained controller in Fig.˜5. In the observability-constrained case, the controller successfully avoids the poor observability region, as can be seen by consistently matching the OpNav measurement and EKF prediction to the true trajectory at all times. An additional test was executed with a [2.5 1 1] ellipsoid using the rotation parameters of Bennu and an approximation from its known mean radius.
Camera snapshots of each case with the detected horizon points highlighted in red can be seen in Fig.˜6. It displays the image history from the spacecraft camera perspective and allows us to make physical inferences about the controller’s performance. For the leftmost case without the observability-constrained control, the spacecraft does not avoid the keep-out cone which causes the lit limb to go out of view in the fourth and fifth images. Thus, the number of detected horizon points significantly decreases, causing the spacecraft to diverge. The limb comes back into view later, but is at a larger distance due to the divergence, making the measurement less effective. The middle case displays the same example but with the observability-constrained control, with which we see that the lit limb is kept within view throughout the trajectory. The rightmost case displays the view of the [2.5 1 1] ellipsoid using observability-constrained control to demonstrate the controller’s consistency in performance across varying shapes of celestial bodies. Although the size of the lit limb varies considerably due to the body’s rotation and viewing the ellipsoidal geometry from different angles, it maintains the lit limb within view by avoiding the poor observability zone and therefore produces more reliable measurements. Both the middle and rightmost cases avoid the divergence and disappearance of lit limb encountered in the leftmost case, supporting the benefit of the observability-constrained controller.
margin: OSIRIS-REx used impulses, inappropriate for this case. Replaced with Hayabusa2’s ion engines. For this trajectory, the average control acceleration was $4.2865\times 10^{-7}\,\textrm{m/s}^{2}$, with a maximum instantaneous acceleration of $4.4163\times 10^{-6}\,\textrm{m/s}^{2}$ and total $\Delta v=0.1687\,\textrm{m/s}$. Using the Hayabusa2 spacecraft’s ion engine system as a reference, the maximum fuel consumption to execute the control used during this trajectory was calculated to be approximately 0.0034 kg over 3 days, which is a reasonable quantity within its fuel reserves. The properties of the Hayabusa2 spacecraft used were an ISP of 3000 seconds and a spacecraft mass of 600 kg [72, 73]. The equation used was $\Delta m=m_{\text{sc}}\cdot\sum\left(u_{\text{vec}}\right)\cdot\Delta t\cdot\frac{1}{\text{Isp}\cdot g_{0}}$ where $m_{\text{sc}}={\color[rgb]{0,0,0}600}\,\text{kg}$ is the spacecraft mass, $u_{\text{vec}}$ is the total acceleration imparted (in $\text{m/s}^{2}$), $\Delta t$ is the time step, $\text{Isp}={\color[rgb]{0,0,0}3000}\,\text{s}$ is the specific impulse, and $g_{0}=9.80665\,\text{m/s}^{2}$ is the standard acceleration due to gravity.
(a) Trajectory (3D view) (b) Trajectory (XY view)
(c) Trajectory (YZ view) (d) Trajectory (XZ view)
(e) Angle to +x-axis (Hill) over time (f) EKF distance error over time
@@
Figure 4: Fig. X Stationkeeping scenario without Observability-constrained Penalty Lyapunov control. True (blue), OpNav measurement (green), EKF-estimated (red) trajectories and keep-out cone (shaded red) are shown in four views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated. +x-axis (Hill) relative angle and EKF distance error plots are included to correlate constraint violation and divergence.
(a) Trajectory (3D view) (b) Trajectory (XY view)
(c) Trajectory (YZ view) (d) Trajectory (XZ view)
(e) Angle to +x-axis (Hill) over time (f) EKF distance error over time
@@
Figure 5: Fig. X Stationkeeping scenario with Observability-constrained Penalty Lyapunov control. True (blue), OpNav measurement (green), EKF-estimated (red) trajectories and keep-out cone (shaded red) are shown in four views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated. +x-axis (Hill) relative angle and EKF distance error plots are included to correlate constraint violation and divergence.
(a) Without observability-constrained control
(b) Constrained (Spherical primary)
(c) Constrained ([2.5 1 1] Ellipsoid primary)
@@
Figure 6: Stationkeeping scenario. Sample image history with detected horizon points from Fig.˜5. Shown are: (a) without observability-constrained control, (b) with observability-constrained control using a spherical primary, and (c) with control using a [2.5 1 1] ellipsoid primary. Images are ordered top–bottom, then left–right.
@@ 5.1.1 Monte Carlo Analysis
To demonstrate the algorithm’s robustness, a Monte Carlo analysis is conducted to analyze its success rate statistically. A pessimistic initial standard deviation is defined for the position error across each direction ($x,y\text{ and }z$ in the Hill frame) of $\sigma=30$ meters, approximately 10% of Bennu’s radius, using the error experienced during the OSIRIS-REx mission as a standard [70, 71]. This is used to apply an initial error to the true initial state to generate the EKF’s initial known state. The Monte Carlo is run for 200 iterations. To start, we test the version without observability constraints to understand the severity of failures in this scenario. As per the results, the Lyapunov controller fails in almost every case without the observability constraint, as it passes through the poor observability region in every test. It succeeds in 8 out of 200 iterations for a success rate of 4%. The EKF and true dynamics diverge completely and lead to the spacecraft either crashing into the small body or traveling out of the system. Now, we compare this with the observability-constrained case. The observability-constrained Lyapunov controller shows better performance and success rate of the spacecraft reaching its desired orbit. It succeeds in 196 out of 200 iterations for a success rate of 98%. This is a major performance improvement and supports the utility of the observability-constrained controller. The ellipsoid case succeeds 161 out of 200 iterations for a success rate of 80.5%. The performance degrades with nonspherical objects but maintains an improvement over the case without an observability constraint. All trajectory Monte Carlo results are shown in Fig.˜7.
Additionally, the evolution of EKF position Norm Error, OpNav measurement Norm Error, angle constraint satisfaction, Lyapunov function value, thrust magnitude over time and a $\Delta v$ consumption histogram across the unconstrained, constrained with sphere and constrained with ellipse test cases are displayed in Fig.˜8. Note that these plots depict error by subtracting truth distance norm from the EKF distance norm or OPNAV measurement norm respectively, leading to many diverging cases for the basic Lyapunov controller that have a negative norm error value due to the spacecraft leaving the system in truth. Likewise, the error bounds are constructed using the EKF position covariance estimate or OpNav analytical covariance. The error magnitude in the unconstrained case is significantly higher than the constrained cases, and has considerable divergence due to poor measurements as the trajectory violates the angle constraint. Meanwhile, the ellipsoid with constraint performs significantly better than the unconstrained case, but experiences slightly more deviation and error than the sphere case with periodic variations in the measurement from the small body’s rotation. In the observability-maintaining cases, the EKF is mostly bounded by the $3\sqrt{\lambda_{max}(P_{\bm{r}})}$ bounds with few outliers, supporting its proper performance. However, the EKF covariance poorly bounds the errors in the basic controller case, corresponding to a deviation between the truth and EKF position estimate, which is concordant with our earlier definition of divergence. Furthermore, the constraints are correctly satisfied by the observability-maintaining Lyapunov controller as the unconstrained scenario enters the dark side cone with an angle lower than $30^{\circ}$ while the constrained cases remain consistently above. For the unconstrained case, The Lyapunov function value starts at a relatively lower value before increasing due to the natural progression of the trajectory, after which the controller attempts to reach the target orbit, in the process of which it encounters divergence since it does not consider the angle constraint. Meanwhile, the constrained cases show a similar pattern with two peaks corresponding to the angle constraint activation, causing the controller to maintain trajectories that maintain observability and thereby minimizing divergence. This is concordant with the continuous thrust profiles depicting behaviour with two peaks corresponding to the angle constraint inflations in the Lyapunov function value. The unconstrained trajectory encounters the maximum thrust constraint due to divergence and therefore has a significantly higher $\Delta v$ usage, while the constrained trajectories have a reasonable thrust distribution.
(a) Without observability-constrained control
(b) Constrained (Spherical primary)
(c) Constrained ([2.5 1 1] Ellipsoid primary)
@@
Figure 7: Stationkeeping scenario. Monte Carlo simulation of 200 true trajectories using (a) basic Lyapunov controller, (b) observability-constrained Lyapunov controller with spherical primary, and (c) with [2.5 1 1] ellipsoid primary. Initial position (green circle), final position (red cross), and orbital direction (black arrows) are indicated.
Figure 8: Stationkeeping scenario. Monte Carlo Simulation of EKF Position Norm Error (Row 1), OpNav Measurement Norm Error (Row 2) and Angle Constraint Satisfaction (Row 3) with 200 Trials using Basic Lyapunov Controller (Column 1), Observability-constrained Lyapunov Controller with Spherical (Column 2) and [2.5 1 1] Ellipsoid Primary (Column 3).
@@ 5.2 Approach and Circularization Scenario
Beyond the orbit maintenance scenario, it is important to demonstrate the effectiveness of this controller in a greater variety of situations, including those starting further away from the small body. Therefore, we test this algorithm in a specific test case where we approach the small body from just within the effective range of horizon-based OpNav determined in Section˜4.1. The controller aims to circularize around the small body and executes maneuvers accordigly. The initial conditions in Cartesian terms are $^{H}\bm{r}=[0,0,-4.5964]^{\top}$ km and $^{H}\bm{v}=[-4.8927,0,2.4464]^{\top}$ mm/s. The same initial covariance is used as in Section˜5.1. Using the tuning guidelines outlined in appendix˜D, the gain found to return the best performance for the controller in this particular test case is:
@@
The penalty function parameters chosen are identical to those from Section˜5.1, except the weight for the cone constraint is increased to $\omega_{3}$ = 100. This leads to better performance in avoiding the poor observability region at the larger scale of this scenario.
@@
We repeat a comparison between the regular (Fig.˜9) and observability-constrained (Fig.˜10) controllers. The figures depict the spacecraft’s true trajectory, OpNav measurement, EKF position estimate and a visualization of the keep-out cone. The result matches the outcome from Section˜5.1 with the observability-constrained controller avoiding the keep-out cone and successfully circularizing around the small body. The regular controller’s true and EKF-estimated trajectories diverge since it enters the keep-out cone and produces faulty measurements. An additional test with the observability-constrained controller was executed with a [2.5 1 1] ellipsoid using the rotation parameters of Bennu and an approximation from its known mean radius.
@@
Camera snapshots of each case with the detected horizon points highlighted in red can be seen in Fig.˜11. The observations are mostly similar to those from Section˜5.1. The difference is that the lit limb is much smaller at the start of the simulation and grows as the spacecraft approaches the small body. The measurement and controller perform well under such variable conditions.
@@
For this trajectory, the average control acceleration was $2.6889\times 10^{-7}\,\textrm{m/s}^{2}$, with a maximum instantaneous acceleration of $6.6203\times 10^{-6}\,\textrm{m/s}^{2}$ and total $\Delta v=0.1234\,\textrm{m/s}$. The maximum fuel consumption to execute the control used during this trajectory was calculated to be approximately 0.0025 kg over 5 days, which is a reasonable quantity within Hayabusa2’s fuel reserves. This calculation was done using the same method from Section˜5.1.
(a) Trajectory (3D view) (b) Trajectory (XY view)
(c) Trajectory (YZ view) (d) Trajectory (XZ view)
(e) Angle to +x-axis (Hill) over time (f) EKF distance error over time
@@
Figure 9: Approach and Circularization scenario without Observability-constrained Penalty Lyapunov control. True (blue), OpNav measurement (green), EKF-estimated (red) trajectories and keep-out cone (shaded red) shown in four views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated. +x-axis (Hill) relative angle and EKF distance error plots included to correlate constraint violation and divergence.margin: Figure Updated with Smoother Trajectory
(a) Trajectory (3D view) (b) Trajectory (XY view)
(c) Trajectory (YZ view) (d) Trajectory (XZ view)
(e) Angle to +x-axis (Hill) over time (f) EKF distance error over time
@@
Figure 10: Approach and Circularization scenario with Observability-constrained Penalty Lyapunov control. True (blue), OpNav measurement (green), EKF-estimated (red) trajectories and keep-out cone (shaded red) with Observability-constrained Penalty Lyapunov control shown in four views. Initial position (green circle), final position (red cross) and orbital direction (black arrows) indicated. +x-axis (Hill) relative angle and EKF distance error plots included to correlate constraint violation and divergence.
(a) Without observability-constrained control
(b) Constrained (Spherical primary)
(c) Constrained ([2.5 1 1] Ellipsoid primary)
@@
Figure 11: Approach and Circularization scenario. Sample image history with detected horizon points from Fig.˜9 for case without observability-constrained control, and Fig.˜10 for cases with observability-constrained control using a spherical primary and a [2.5 1 1] ellipsoid primary. Images are ordered top–bottom, then left–right.